A 26 mal solution containing 2.0 M HClO3 is mixed with a 40mL solution containing 0.5 M NaOH. What is the pH of the resulting mixture ?


1
Expert's answer
2021-06-14T13:36:23-0400

Reaction:

"HClO_3" is a strong aciid

"HClO_3+NaOH=NaClO_3+H_2O"


"\nu_0(NaOH)=C(NaOH)\cdot V(NaOH)=0.02 mol""\nu_0(HClO_3)=C(HClO_3)\cdot V(HClO_3)=0.052 mol"

"\nu(NaOH)=0mol"

"\nu(HClO_3)=(0.052-0.02) mol=0.032 mol"


Total volume:

"V = V_1+V_2=66ml=.066L"


"C(HClO_3)=[H^+]=\nu(HClO_3)/V=0.485M"

"pH=-lg[H^+]=0.314"